# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ V Numerical Results
To validate the aforementioned discussion on the roots of the characteristic function, Fig. 2(b) presents the effect of gradient staleness on the convergence rate of SGD with gradient staleness. To avoid the impact of gradient noise on SGD, we set $\sigma=0$. Also, the learning rate is $\eta=0.01$, and we adopt the same objective function as in Fig. 2(a). We observe that, as the discrete-time gradient staleness increases from 0 to 36, an increased degree of gradient staleness leads to faster convergence. While as the gradient staleness further increases, i.e., when $\eta\breve{\tau}>{1}/{e}$, SGD oscillates, which leads to slower converges. Further, it can be seen that when $\breve{\tau}\sim N(36,5^{2})$, SGD nearly has the nearly same converges rate as when the discrete-time gradient staleness is a constant given by $\mathbb{E}\{\breve{\tau}\}$.
@@
To reveal the optimal choice of the learning rate in multi-dimensional optimization, Fig. 3(b) is presented. In particular, $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}\boldsymbol{x}$ in which $\boldsymbol{V}=\text{diag}(v,1)$. Notably, $\left\|\boldsymbol{V}\right\|_{2}=\max\{v,1\}$. Given the discrete-time gradient staleness and $v$, we numerically find the optimal (fixed) step size, i.e., the step-size that leads to the minimum loss after $1\times 10^{5}$ iterations, by exhaustive search. For example, it is seen that $\eta^{*}=\frac{1}{e\breve{\tau}}$ when $v=1$, which agrees with Eq. (18). Notably, we observe that it always holds that $\eta^{*}\in\left[1/{(e\left\|\boldsymbol{V}\right\|_{2})},{\pi}/{(2\left\|\boldsymbol{V}\right\|_{2})}\right)$, which is consistent with Eq. (19) in Subsection III-C. This is due to the fact that, when $\eta>{\pi}/{(2\left\|\boldsymbol{V}\right\|_{2})}$, in the eigenvector direction which the largest eigenvalue, SGD diverges. While SGD converges slower along each eigenvector direction by further reducing the learning rate when $\eta<{1}/{(e\left\|\boldsymbol{V}\right\|_{2})}$.
@@
In the subsequent, we verify our differential equation-based performance analysis on the effect of the gradient noise in Fig. 3(a). To avoid the impact of gradient staleness on the convergence of SGD, we study $K$-synchronized SGD. The objective functions are $f_{i}(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}_{i}\boldsymbol{x}$, in which $\boldsymbol{V}_{1}=\text{diag}(1,1,1)$, $\boldsymbol{V}_{2}=\text{diag}(1,1,3)$, and $\boldsymbol{V}_{3}=\text{diag}(1,2,3)$. Also, the variance of the additive Gaussian noise is inversely proportional to the square root of $K$, in which $K$ is the number of workers. It is seen that in quadratic optimization, the number of iterations for an SGD algorithm is mainly determined by the smallest eigenvalue. Further, it can be observed that the average iterations for the SGD algorithm to meet the stopping criterion stays nearly unchanged despite the increases in the number of users with the objective function $f_{3}(\boldsymbol{x})$. This is due to the fact that the average first hitting time of a one-dimensional OU-process remains unchanged with respect to the variance of the additive Gaussian noise as suggested by Lemma 1. Further, it is seen that the training time slightly decreases with the increase of $K$ with $\boldsymbol{V}_{1}$, which fits well with Eq. (13). This is due to the fact the decrease in the noise variance leads to a decrease in the variance of the first hitting times along each eigenvector. In addition, we observe that, the differential-equation based approach presents an accurate performance analysis with a quadratic function. Also, it should be noted that the increase of $K$ reduces the variance of the training time.
@@
In the following simulations, we investigate the effect of gradient noise and staleness and verify our performance analysis based on empirical results on the MNIST handwritten digit database. First, we present the effect of the number of users on asynchronous learning with a fixed learning rate. In particular, the learning rate is $\eta=0.01$. The computational delay of each worker follows an i.i.d. Gaussian distribution $\mathcal{N}(10,1^{2})$ and the communication delay is assumed to be negligible.
(a) Number of iterations required for synchronous SGD with $K$ workers to converge.
(b) The optimal learning rate for asynchronous SGD.
Figure 3: The impact of the gradient noise and the choice of learning rate.
@@
Fig. 5(a) presents the performance of asynchronous learning with respect to wall-clock time on the MNIST dataset with a convolutional neural network (CNN). First, similar to the solvable cases, the performance of asynchronous learning first increases with the increase of $K$ (or step staleness), then decreases. It is seen that when there are fewer workers, for example, $K=5$, asynchronous SGD converges slowly. This is due to the fact that with small $K$, the number of gradients collected by the global model is also limited. Further, when $K=20$, the asynchronous SGD algorithm achieves nearly the best performance. Another observation that can be made from Fig. 5(a) is that when $K\geq 20$, the performance of asynchronous SGD deteriorates, which diverges when $K\geq 40$. The reason for this is that gradients staleness becomes significant as $K$ increases, which leads to the divergence of asynchronous SGD. Note that the discrete-time gradient staleness is of significance in our analysis, which is defined as the number of global model updates between two consecutive gradient updates for the same user. This is different from the continuous-time gradient staleness with respect to the wall-clock time. These results demonstrate that, too many workers in asynchronous learning can considerably deteriorate the efficiency of the algorithms. [^16] In the simulation, we have also presented a rough estimation of the norm-2 of the Hessian matrix $\boldsymbol{H}$ of the employed CNN, which indicates that $\left\|\boldsymbol{H}\right\|_{2}$ falls in the interval $(3,5)$ during the first several iterations. When $K=20$, asynchronous SGD almost achieves the optimal performance. In this case, we have $\mathbb{E}\{\breve{\tau}\}\approx 20$. Hence, we have $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}\in(0.6,1.0)$. Note that, for the solvable cases, a necessary condition for asynchronous SGD to achieve optimal performance is $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}\in({1}/{e},{\pi}/{2})$. Also, asynchronous SGD diverges when $\mathbb{E}\{\breve{\tau}\}\geq 40$, which agrees with the argument in Eq. (19) for solvable cases in which asynchronous SGD diverges when $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}>\pi/2$. This suggests the potential that the results for the solvable cases can be extended to general non-convex cases.
@@
(a) Time-performance of asynchronous SGD versus the number of clients with a fixed learning rate.
@@
(b) Time-performance of asynchronous SGD versus the number of clients with a worker number-aware learning rate.
Figure 5: Asynchronous SGD with gradient staleness on the MNIST dataset.
@@
With a fixed step-size, the performance of asynchronous SGD is shown to deteriorate with a large number of workers, or equivalently, with a large degree of gradient staleness. Next, we aim to evaluate the performance of asynchronous SGD with a worker number-aware step-size. In particular, we adopt the same setup as the experiment presented in Fig. 5(a). Fig. 5(b) presents the performance of asynchronous SGD with worker number-aware step size. In particular, the step size is $\eta(K)=\frac{\eta_{0}}{K}$, wherein $K$ is the number of workers and $\eta_{0}=0.2$ is a constant. In this case, it is seen that the asynchronous SGD algorithm converges with a worker number-aware step size. However, a careful reader may also notice that the performance of the SGD algorithm does not significantly improve with the increase of $K$ in terms of wall-clock time. This is due to the fact that the positive effect of the increased number of gradients as $K$ increases is offset by the step size inversely proportional to the number of users. While asynchronous learning can outperform synchronous learning with respect to the wall-clock time with a relatively small $K$. This observation aligns with the literature, such as [14].
@@ VI Conclusion
In this work, we have presented performance analysis and protocol design criteria for distributed SGD with gradient staleness through an SDDE-based approach. To deal with non-exponentially distributed computation time, a Poisson approximation for the superposition of independent renewal processes has been adopted, based on which we have presented the run-time and staleness analysis of asynchronous SGD. In this way, we have bridged the behavior of the asynchronous SGD algorithm with gradient staleness and the solution of certain SDDE. We have derived the convergence condition of distributed SGD by analyzing its characteristic roots, which allows us to optimize scheduling policies for asynchronous distributed SGD. In particular, the characteristic roots have been shown to be closely related to the learning rate, step staleness, and the eigenvalues of the objective function’s Hessian matrix. It has been shown that the presence of a small gradient could slightly accelerate the convergence of the SGD, while a large degree of gradient staleness leads to its convergence. It has also been shown that, regardless of the specific distribution of computation time, the expectation of the step staleness in asynchronous SGD without gradient dropout is determined only by the number of workers and the group size. With limited bandwidth, excessive workers can lead to a large communication delay due to network congestion, degrading the performance of SGD. Further, we have also extended the derivation of the characteristic roots of the case where the gradient staleness in SDDE follows a uniform distribution, which arises in event-triggered SGD. In practice, for a given learning rate, it has been shown that there exists an optimal number of workers in the asynchronous SGD tasks, for the fact that a large number of workers leads to a greater degree of staleness which may lead to divergence of the SGD algorithm. Numerical simulation results have been presented to demonstrate the potential of our SDDE framework, even in complex learning tasks with non-convex objective functions.
It is noteworthy that the step staleness or discrete-time staleness, rather than the gradient delay, plays a crucial role in determining the convergence rate of asynchronous SGD. While asynchronous learning has been demonstrated to outperform synchronous learning with respect to the wall-clock time with a small number of users, the performance of asynchronous learning does not significantly improve with additional workers, mainly due to the enlarged gradient staleness. In this paper, we have mainly focused on the solvable cases, but we have also provided quite suggestive and analytical discussion. Important potential directions for extending our works involve extending the analytical SDDE-based framework of asynchronous SGD with gradient staleness to nonconvex objective function, and performance analysis of asynchronous SGD with a staleness-aware learning rate or involving a wireless channel.
Appendix A Proof of Corollary 4

### v2

@@ V Numerical Results
To validate the aforementioned discussion on the roots of the characteristic function, Fig. 2(b) presents the effect of gradient staleness on the convergence rate of SGD with gradient staleness. To avoid the impact of gradient noise on SGD, we set $\sigma=0$. Also, the learning rate is $\eta=0.01$, and we adopt the same objective function as in Fig. 2(a). We observe that, as the discrete-time gradient staleness increases from 0 to 36, an increased degree of gradient staleness leads to faster convergence. As the gradient staleness further increases, i.e., when $\eta\breve{\tau}>{1}/{e}$, SGD oscillates, which leads to slower convergence. Further, it can be seen that when $\breve{\tau}\sim N(36,5^{2})$, SGD nearly has the same convergence rate as when the discrete-time gradient staleness is a constant given by $\mathbb{E}\{\breve{\tau}\}$.
@@
To reveal the optimal choice of the learning rate in multi-dimensional optimization, Fig. 3(b) is presented. In particular, $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}\boldsymbol{x}$ in which $\boldsymbol{V}=\text{diag}(v,1)$. Notably, $\left\|\boldsymbol{V}\right\|_{2}=\max\{v,1\}$. Given the discrete-time gradient staleness and $v$, we numerically find the optimal (fixed) step size, i.e., the step-size that leads to the minimum loss after $1\times 10^{5}$ iterations, by exhaustive search. For example, we observe that $\eta^{*}=\frac{1}{e\breve{\tau}}$ when $v=1$, which agrees with Eq. (29). Notably, we observe that it always holds that $\eta^{*}\in\left[1/{(e\left\|\boldsymbol{V}\right\|_{2})},{\pi}/{(2\left\|\boldsymbol{V}\right\|_{2})}\right)$, which is consistent with Eq. (30) in Subsection III-C. This is due to the fact that, when $\eta>{\pi}/{(2\left\|\boldsymbol{V}\right\|_{2})}$, in the eigenvector direction which the largest eigenvalue, SGD diverges. Additionally, SGD converges slower along each eigenvector direction by further reducing the learning rate when $\eta<{1}/{(e\left\|\boldsymbol{V}\right\|_{2})}$.
(a) Number of iterations required for synchronous SGD to converge.
(b) The optimal learning rate for ASGD.
Figure 3: The impact of the gradient noise and learning rate on the convergence of the gradient descent algorithms.
@@
In the subsequent, we verify our differential equation-based performance analysis on the effect of the gradient noise in Fig. 3(a). To avoid the impact of gradient staleness on the convergence of SGD, we study $K$-synchronized SGD. The objective functions are $f_{i}(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}_{i}\boldsymbol{x}$, in which $\boldsymbol{V}_{1}=\text{diag}(1,1,1)$, $\boldsymbol{V}_{2}=\text{diag}(1,1,3)$, and $\boldsymbol{V}_{3}=\text{diag}(1,2,3)$. Also, the variance of the additive Gaussian noise is inversely proportional to the square root of $K$, in which $K$ is the number of workers. It is seen that in quadratic optimization, the number of iterations for an SGD algorithm is mainly determined by the smallest eigenvalue. Further, it can be observed that the average iterations for the SGD algorithm to meet the stopping criterion stays nearly unchanged despite the increase in the number of workers with the objective function $f_{3}(\boldsymbol{x})$. This is due to the fact that the average first hitting time of a one-dimensional OU-process remains unchanged with respect to the variance of the additive Gaussian noise as suggested by Lemma 1. Further, it is seen that the training time slightly decreases with the increase of $K$ with $\boldsymbol{V}_{1}$, which fits well with Eq. (23). This is due to the fact that the decrease in the noise variance leads to a decrease in the variance of the first hitting times along each eigenvector. In addition, we observe that, the differential-equation based approach presents an accurate performance analysis with a quadratic function. Also, it should be noted that the increase of $K$ reduces the variance of the training time.
@@
In the following simulations, we investigate the effect of gradient noise and staleness and verify our performance analysis based on empirical results on the MNIST handwritten digit database. First, we present the effect of the number of workers on asynchronous learning with a fixed learning rate. In particular, the learning rate is $\eta=0.01$. The computational delay of each worker follows an i.i.d. Gaussian distribution $\mathcal{N}(10,1^{2})$ and the communication delay is assumed to be negligible. In this case, we employ a convolutional neural network (CNN). More specifically, the neural network consists of two convolutional layers followed by a fully connected layer. The first conventional layer takes a single-channel input and applies 16 filters with a kernel size of 5. It has a stride of 1 and uses padding of 2 to keep the output size the same. After the convolution, it applies a ReLU activation function and then performs max pooling with a kernel size of 2. The second layer takes the output from the first layer and applies 32 filters with the same kernel size and parameters. It also uses ReLU activation and max pooling to further refine the features. The output from the second convolutional layer is flattened into a one-dimensional vector and fed into a fully connected layer that outputs 10 values, corresponding to digits 0 to 9. Additionally, the Hessian matrix is evaluated after 10 iterations from the start of the training process and the loss function selected for training is the cross-entropy loss.
@@
Fig. 5(a) presents the performance of asynchronous learning with respect to wall-clock time on the MNIST dataset with a CNN. First, similar to the solvable cases, the performance of asynchronous learning first increases with the increase of $K$ (or step staleness), then decreases. It is seen that when there are fewer workers, for example, $K=5$, ASGD converges slowly. This is due to the fact that with small $K$, the number of gradients collected by the global model is also limited. Further, when $K=20$, the ASGD algorithm achieves nearly the best performance. Another observation that can be made from Fig. 5(a) is that when $K\geq 20$, the performance of ASGD deteriorates, which diverges when $K\geq 40$. The reason for this is that gradient staleness becomes significant as $K$ increases, which leads to the divergence of ASGD. Note that the discrete-time gradient staleness is of significance in our analysis, which is defined as the number of global model updates between two consecutive gradient updates for the same worker. This is different from the continuous-time gradient staleness with respect to the wall-clock time. These results demonstrate that, too many workers in asynchronous learning can considerably deteriorate the efficiency of the algorithms. [^12] In the simulation, we have also presented a rough estimation of the norm-2 of the Hessian matrix $\boldsymbol{H}$ of the employed CNN, which indicates that $\left\|\boldsymbol{H}\right\|_{2}$ falls in the interval $(3,5)$ during the first several iterations. When $K=20$, ASGD almost achieves the optimal performance. In this case, we have $\mathbb{E}\{\breve{\tau}\}\approx 20$. Hence, we have $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}\in(0.6,1.0)$. Note that, for the solvable cases, a necessary condition for ASGD to achieve optimal performance is $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}\in({1}/{e},{\pi}/{2})$. Also, ASGD diverges when $\mathbb{E}\{\breve{\tau}\}\geq 40$, which agrees with the argument in Eq. (30) for solvable cases in which ASGD diverges when $\eta\mathbb{E}\{\breve{\tau}\}\left\|\boldsymbol{H}\right\|_{2}>\pi/2$. This suggests the potential that the results for the solvable cases can be extended to general non-convex cases.
@@
(a) Time-performance of ASGD versus the number of clients with a fixed learning rate.
@@
(b) Time-performance of ASGD versus the number of clients with a worker number-aware learning rate.
Figure 5: ASGD with gradient staleness on the MNIST dataset.
@@
With a fixed step-size, the performance of ASGD is shown to deteriorate with a large number of workers, or equivalently, with a large degree of gradient staleness. Next, we aim to evaluate the performance of ASGD with a worker number-aware step-size. In particular, we adopt the same setup as the experiment presented in Fig. 5(a). Fig. 5(b) presents the performance of ASGD with worker number-aware step size. In particular, the step size is $\eta(K)=\frac{\eta_{0}}{K}$, wherein $K$ is the number of workers and $\eta_{0}=0.2$ is a constant. In this case, it is seen that the ASGD algorithm converges with a worker number-aware step size. However, a careful reader may also notice that the performance of the SGD algorithm does not significantly improve with the increase of $K$ in terms of wall-clock time. This is due to the fact that the positive effect of the increased number of gradients as $K$ increases is offset by the step size inversely proportional to the number of workers. While asynchronous learning can outperform synchronous learning with respect to the wall-clock time with a relatively small $K$. This observation aligns with the literature, such as [16].
@@ VI Conclusion
In this work, we have presented performance analysis and protocol design criteria for asynchronous SGD with gradient staleness through an SDDE-based approach. To deal with non-exponentially distributed computation time, a Poisson approximation for the superposition of independent renewal processes has been adopted, based on which we have presented the run-time and staleness analysis of ASGD. In this way, we have bridged the behavior of the ASGD algorithm with gradient staleness and the solution of certain SDDE. We have derived the convergence condition of distributed SGD by analyzing its characteristic roots, which allows us to optimize scheduling policies for asynchronous distributed SGD. In particular, the characteristic roots have been shown to be closely related to the learning rate, step staleness, and the eigenvalues of the objective function’s Hessian matrix. This work has also demonstrated that ASGD has a higher error floor compared to synchronous SGD. However, this issue can be mitigated by reducing the level of asynchrony as ASGD approaches the local minimum. It has been shown that the presence of a small gradient could slightly accelerate the convergence of the SGD, while a large degree of gradient staleness leads to its convergence. It has also been shown that, regardless of the specific distribution of computation time, the expectation of the step staleness in ASGD without gradient dropout is determined only by the number of workers and the group size. With limited bandwidth, excessive workers can lead to a large communication delay due to network congestion, degrading the performance of SGD. Further, we have also extended the derivation of the characteristic roots of the case where the gradient staleness in SDDE follows a uniform distribution, which arises in event-triggered SGD. In practice, for a given learning rate, it has been shown that there exists an optimal number of workers in the ASGD tasks, for the fact that a large number of workers leads to a greater degree of staleness which may lead to divergence of the SGD algorithm. Numerical simulation results have been presented to demonstrate the potential of our SDDE framework, even in complex learning tasks with non-convex objective functions.
ASGD benefits from a higher gradient arrival rate compared to synchronous SGD. It is noteworthy that the step staleness or discrete-time staleness, rather than the gradient delay, plays a crucial role in determining the convergence rate of ASGD. It has been demonstrated that, through careful selection of the number of workers or by designing appropriate protocols, step staleness can be effectively controlled. Also, it is possible to leverage stale gradients to accelerate the convergence of ASGD. Important theoretical potential directions for extending our SDDE-based works involve further analytical results with a non-convex objective function, staleness-aware learning rates, ASGD variants, and non-Gaussian gradient noise. Important practical future directions include developing adaptive learning rates and designing protocols that leverage gradient staleness to accelerate the convergence of ASGD.
Acknowledgement
The authors would like to thank the editor and anonymous reviewers for their professional and constructive comments which helped to improve the quality of this paper.
Appendix A Proof of Proposition 2
Without loss of generality, we consider the hitting time to $\delta>0$ rather than $-\delta$. Hence, in this proof, let $x(0)>0$ such that the OU process first reaches $\delta>0$. Also, the subscript $i$, which denotes the $i$-th OU process is omitted in this proof for brevity. First, we aim to show that the probability of the OU process leaving the interval $(-\infty,\delta)$ after it enters this interval can be bounded by a small positive number $\epsilon>0$.
Let $k_{0}$ be the solution to $x(k_{0})=\delta$. Then it holds that $\mathbb{E}\{x(k_{0}+k\eta)\}=x_{0}\exp(-{v}k\eta)$ and $\text{Var}\{x(k_{0}+k\eta)\}=\frac{\sigma^{2}}{2{v}}(1-\exp(-2{v}k\eta))$ where $x(k_{0}+k\eta)$ is a Gaussian distributed random variable. Therefore, we have $\Pr(x(k_{0}+k\eta)>\delta)\overset{(a)}{\leq}Q\left(\frac{\sqrt{2{v}}\delta}{\sigma}(1-\exp(-{v}k\eta)\right)$ where $(a)$ follows from the fact that $1-\exp(-2{v}k\eta)\leq 1$ with positive ${v},k,$ and $\eta$. Note that in practice, the evaluation of the value of the objective variable is only conducted at a few discrete time points. Let $E$ denote the event that $x(k_{0}+k\eta)>\delta$ for any $k\in(k_{0}+k_{l},k_{0}+k_{r})$, where $k\in\mathbb{N}^{+}$, $10<\epsilon,\beta<$ 1, $k_{l}=\frac{-\log(\beta)}{{v}\eta}$, and $k_{r}=\frac{\epsilon}{\eta\exp\left(-\frac{(1-\beta)^{2}\delta^{2}{v}}{\sigma^{2}}\right)}$. In the following, we aim to demonstrate the fact that
$P(E)\leq\epsilon.$ (53)
Clearly, for $k>k_{l}$, it holds that $\Pr(x(k_{0}+k\eta)>\delta)\leq\exp\left(-\frac{{v}\delta^{2}}{\sigma^{2}}(1-\beta)^{2}\right)$. Therefore,
$P(E)\leq k_{r}\eta\exp\left(-\frac{{v}\delta^{2}}{\sigma^{2}}(1-\beta)^{2}\right)\leq\epsilon.$ (54)
In this way, it can be seen that once the OU process enters the interval $(-\infty,\delta)$, the probability that the OU process leaves the interval is quite small. Further, we have $\Pr\{{H}\geq t\}=\Pr\{{H}\geq t|x(t)\geq\delta\}\Pr\{x(t)\geq\delta\}+\Pr\{{H}\geq t|x(t)<\delta\}\Pr\{x(t)<\delta\}.$ Clearly, $\Pr\{{H}\geq t|x(t)<\delta\}=0$. Next, it holds that $\Pr\{{H}\geq t|x(t)\geq\delta\}+\Pr\{{H}<t|x(t)\geq\delta\}=1$. By Eq. (53), we make the approximation that $Pr\{{H}<t|x(t)\geq\delta\}\approx 0$, i.e., $\Pr\{{H}\geq t\}\approx\Pr\{x(t)\geq\delta\}.$
Therefore, $\Pr\{x(t)\geq\delta\}=Q\left(\frac{x_{0}\exp(-{v}t)-\delta}{\sigma_{o}(t)}\right)$, where $Q(\cdot)$ is the $Q$-function and $\sigma_{o}(t)=\frac{\sigma}{\sqrt{2{v}}}\sqrt{1-\exp{(-2{v}t)}}$. Let $p_{H}(t)=-\frac{\rm d}{\rm dt}\Pr\{x(t)\geq\delta\}$, which is the p.d.f. of the first hitting time. In this case, we have
$p_{H}(t)=\exp\left(-{\left(\frac{x_{0}e^{-{v}t}-\delta}{\sqrt{2}\sigma_{o}(t)}\right)^{2}}\right)\times$ (55)
$\left[\frac{-{v}x_{0}\exp(-{v}t)}{\sigma_{o}(t)}-{\frac{(x_{0}\exp(-{v}t)-\delta)\frac{\sigma{v}\exp(-2{v}t)}{\sqrt{2{v}(1-\exp(-2{v}t))}}}{(\sigma_{o}(t))^{2}}}\right].$
Here, we are interested in the expression of Eq. (55) around $t=\frac{1}{{v}}\log\frac{x_{0}}{\delta}$. First, note that for $t=t_{0}+\Delta t$, $G(t)=x_{0}\exp(-{v}t)-\delta$ is on the order of $\mathcal{O}(\Delta t)$. This is due to the fact that $x_{0}\exp(-{v}t)-\delta=\delta(\exp(-{v}\Delta t)-1)$. Further, Therefore, $x_{0}\exp(-{v}t)-\delta\approx-\delta{v}\Delta t$ by the first order Taylor expansion. Hence, it holds that $\lim_{\Delta_{t}\to 0}G_{2}(t_{0}+\Delta t)=0$. By ignoring the contribution of $G(t)$ as $\Delta t\to 0$, we have
$p_{H}(t_{0}+\Delta t)=\frac{{v}\delta\exp(-{v}\Delta t)}{\sigma_{o}(t_{0}+\Delta t)}\exp\left({-\frac{{v}\delta(\Delta t)^{2}}{2\sigma^{2}_{o}(t_{0}+\Delta t)}}\right)+\mathcal{O}(\Delta t).$ (56)
By approximating $\exp(-v\Delta t)\approx 1$ and $\sigma_{o}(t_{0}+\Delta t)\approx\sigma_{o}(t_{0})$ for small $\Delta t$, where the approximation error is on the order of $\mathcal{O}(\Delta t)$, we find that
$p_{H}(t_{0}+\Delta t)=\frac{1}{\tilde{\sigma}\sqrt{2\pi}}\exp\left(-\frac{(\Delta t)^{2}}{2\tilde{\sigma}^{2}}\right)+\mathcal{O}(\Delta t),$ (57)
where $\tilde{\sigma}=\sqrt{\frac{\sigma^{2}}{2v_{i}^{3}\delta^{2}}\left(1-\frac{1}{\alpha_{i}^{2}}\right)}$.In practice, $\mathcal{O}(\Delta t)$ is on the same order of $\mathcal{O}(\sigma)$, which suggests that this approximation is accurate, especially with a small gradient noise, i.e., small $\sigma$. Thus, we obtain Eq. (21).
Appendix B A Bound on the Probability of the Event $H_{\max}\neq\tilde{H}$.
A careful reader may observe that the $H_{\max}$ is not necessarily equal to $\tilde{H}$. More specifically, define $i=\arg\max H_{j}$. where $H_{j}$ is given by $H_{j}=\inf\{t>0:|x_{j}(t)-x_{j}^{*}|<\delta\}$, as defined in Eq. (19). Hence, if $|x_{j}(H_{\max})|\leq\delta$ for all $j\neq i$, then $H_{\max}$ and $\tilde{H}$ are identical. Otherwise, if $|x_{j}(H_{\max})|>\delta$ for any $j\neq i$, then we have $H_{\max}\neq\tilde{H}$. Let $L_{j}=|H_{j}-H_{\max}|$, representing the time difference between $H_{j}$ and $H_{\max}$. Next, we would like to demonstrate the probability of the event $|x_{j}(H_{\max})|>\delta$ decays exponentially with $L_{j}$. Specifically, we can establish that
$\text{Pr}(|x_{j}(H_{\max})|>\delta)\overset{(a)}{\leq}2\exp\left(-\frac{v_{j}^{3}\delta^{2}L_{j}^{2}}{\sigma^{2}}\right),$ (58)
where $Q(\cdot)$ is the $Q$-function and the inequality $(a)$ holds by the Chernoff bound of the $Q$-function, i.e., $Q(s)\leq e^{-\frac{s^{2}}{2}}$, $s>0$. Therefore, we conclude that $\text{Pr}(|x_{j}(H_{\max})|>\delta)$ is on the order of $\mathcal{O}(e^{-L_{j}^{2}})$, and hence it becomes negligible for large values of $L_{j}$.
Appendix C Proof for Corollary 6
Let $\omega=\zeta\lambda$. It holds that $\zeta=\frac{\omega^{2}}{e^{-\omega}-1}$ by Eq. (41). Let us define $g(\omega)=\frac{\omega^{2}}{e^{-\omega}-1}$. Therefore, the derivative of $g(\omega)$ with respect to $\omega$ is given by
$g^{\prime}(\omega)=\frac{2\omega(e^{-\omega}-1)+\omega^{2}e^{-\omega}}{(e^{-\omega}-1)^{2}}.$ (59)